# MolecularGPT: Open Large Language Model (LLM) for Few-Shot Molecular Property Prediction

### v1

@@ Abstract
Molecular property prediction (MPP) is a fundamental and crucial task in drug discovery. However, prior methods are limited by the requirement for a large number of labeled molecules and their restricted ability to generalize for unseen and new tasks, both of which are essential for real-world applications. To address these challenges, we present MolecularGPT for few-shot MPP. From a perspective on instruction tuning, we fine-tune large language models (LLMs) based on curated molecular instructions spanning over 1000 property prediction tasks. This enables building a versatile and specialized LLM that can be adapted to novel MPP tasks without any fine-tuning through zero- and few-shot in-context learning (ICL). MolecularGPT exhibits competitive in-context reasoning capabilities across 10 downstream evaluation datasets, setting new benchmarks for few-shot molecular prediction tasks. More importantly, with just two-shot examples, MolecularGPT can outperform standard supervised graph neural network methods on 4 out of 7 datasets. It also excels state-of-the-art LLM baselines by up to 16.6% increase on classification accuracy and decrease of 199.17 on regression metrics (e.g., RMSE) under zero-shot. This study demonstrates the potential of LLMs as effective few-shot molecular property predictors. The code is available at https://github.com/NYUSHCS/MolecularGPT.
@@ 1 Introduction
  • We extensively experimented on 10 molecular property benchmarks across different scales and tasks to validate the effectiveness of MoelcularGPT. Our empirical results demonstrate that MoelcularGPT outperforms the leading LLM baselines (e.g., GIMLET, LLaMA-7b (Touvron et al., 2023), and LLaMA-13B (Touvron et al., 2023)), with up to an average 16.6% improvement across all classification tasks. Additionally, with just two-shot examples, MolecularGPT surpass standard supervised GNN methods on 4 out of 7 datasets, setting new benchmarks for few-shot molecular property tasks.
@@ 2 Related work
GNNs-based MMP
GNNs (Velickovic et al., 2017; Xu et al., 2019; Kipf and Welling, 2017; Stärk et al., 2022) perform MPP tasks by constructing models between molecular graphs and properties. Though have achieved great success (Gilmer et al., 2017; Hamilton et al., 2017; Li et al., 2022; Liu et al., 2022; Tan et al., 2023), these supervised models solely utilize structure information, neglecting the wealthy knowledge contained in texts derived from wet lab experiments. More importantly, they are implicitly trained for each task without explicit natural language instructions, which can not directly generalize to new tasks.
Pretrain-finetune based molecular language models
To utilize the chemical knowledge in texts, molecular language models (Liu et al., 2023b; Edwards et al., 2022; Pei et al., 2023; Zhang et al., 2023; Liu et al., 2023d) aim to integrate natural language and molecular representations for joint reasoning. These models (Su et al., 2022; Zeng et al., 2022; Liu et al., 2023c, 2024; Li et al., 2024) involve two stages: pre-training and fine-tuning. The pre-training phase primarily focuses on learning molecular representations and their associated textual descriptions through masked language modeling, contrastive learning or next token prediction. However, they still require fine-tuning on particular MPP downstream tasks, thereby limiting their generalization abilities to new tasks.
Instruction tuning based molecular language models
To address this, recent efforts in molecular language modeling (Fang et al., 2023; Zhao et al., 2024) aim to explicitly align molecular graphs with their properties through instruction tuning Longpre et al. (2023). For instance, GIMLET (Zhao et al., 2024) integrates molecular graphs with instruction languages for fine-tuning LLMs. GIMLET achieves effective zero-shot ICL for new tasks but lacks few-shot ICL capability due to its generalized position embedding and decoupled attention designs. Mol-Instructions (Fang et al., 2023) is a close work to us, but it fine-tunes LLMs with only three properties tasks and neglects inter-molecular correlations, significantly limiting its zero-shot and few-shot ICL performances. In contrast, we curate a diverse instruction set covering 1000 property tasks and introduce structure-aware few-shot instructions to significantly enhance the zero-shot and few-shot reasoning capabilities of LLMs in MPP tasks. More details about these property tasks can be found in Appendix A.1
@@ 3.1 SMILES-based Molecular Instruction Tuning: A Unified Step
To improve the generalization capability of fine-tuned LLM for MPP tasks (C1), prior models often utilize GNNs (Seidl et al., 2023) or graph transformer (Zhao et al., 2024) as encoders to map molecular graphs into hidden representations. When a graph encoder is well-trained, it can be used to map molecules in $D$ or $D_{\text{test}}$ into a shared hidden space, providing a unified hidden expression. However, as discussed above, this assumption may not hold in practice, as training a unified graph encoder for cross-domain molecules still remains an open-question (Liu et al., 2023a).
To address this, we aim to employ the well-known graph-derived linear strings (Weininger, 1988; Krenn et al., 2020) of molecular graphs, such as SMILES (Weininger, 1988), for instruction tuning. Unlike GNN encoders, SMILES translates molecules’ chemical structure into a string of atomic symbols and chemical bonds (single, double, or triple) based on a set of rules (Qian et al., 2023). This precise translation not only accounts for the graph structure within each molecular graph, but also generalizes readily to arbitrary molecular graphs, providing a universal expression foundation for different types of molecules. Following standard instruction tuning protocol (Christofidellis et al., 2023; Fang et al., 2023; Zhang et al., 2023; Liu et al., 2024; Li et al., 2024), the molecular instruction set $S_{D}$ can be generated by the following prompt template $T=\{Q,I,R\}$ based on $D$, regarding as a zero-shot instruction template.
@@ 3.3 Hybrid Molecular Instruction Tuning: Better Few-Shot Learner
This observation motivates us to introduce a hybrid instruction set $S_{D}^{h}$, combining the strengths of both the zero-shot instruction template $T$ and the few-shot instruction template $T_{shot}$. Specifically, $S_{D}^{h}$ is derived from a combination of 0, 1, 2, 3, and 4-shot instruction templates. In Fig. 2, we can see that our hybrid instruction tuned models, 0&4-shot_tuning and 0-4-shot_tuning, consistently outperforms others in both zero-shot and one-shot scenarios. Further details can be found in Section 4.3.
@@ 4 Experiment
Table 1: Performance over Bio-activity, Toxicity, and Pharmacokinetic classification tasks. Highlights are the first, second, and third best results of zero- and few-shot performances. In supervised finetuned models, we also mark the highest and lowest results.
Method Model Size Type BACE HIV MUV Avg.bio Tox21 ToxCast Avg.tox BBBP CYP450 Avg.pha
XVPLM 110M 0.5126 0.6120 0.6172 0.5806 0.4917 0.5096 0.5007 0.6020 0.5922 0.5971
MoMu 113M 0.6656 0.5026 0.6051 0.5911 0.5757 0.5238 0.5498 0.4981 0.5798 0.5390
Galactica-125M 125M 0-Shot 0.4451 0.3671 0.4986 0.4369 0.4964 0.5106 0.5035 0.6052 0.5369 0.5711
Galactica-1.3B 1.3B 0.5648 0.3385 0.5715 0.4916 0.4946 0.5123 0.5035 0.5394 0.4686 0.5040
GIMLET 64M 0.6957 0.6624 0.6439 0.6673 0.6119 0.5904 0.6011 0.5939 0.7125 0.6532
0-shot 0.4911 0.6060 0.5554 0.5508 0.5481 0.4693 0.5087 0.3671 0.4198 0.3935
1-shot 0.4911 0.6060 0.5554 0.5508 0.5481 0.4954 0.5218 0.3671 0.4198 0.3935
LLaMA2-chat-7B 7B 2-shot 0.6930 0.6587 0.5085 0.6201 0.6052 0.5010 0.5531 0.5459 0.5807 0.5633
4-shot 0.7685 0.6781 0.4685 0.6384 0.6199 0.5025 0.5612 0.5423 0.6092 0.5758
6-shot 0.7180 0.7058 0.5133 0.6457 0.6334 0.5228 0.5781 0.5161 0.6145 0.5653
0-shot 0.6561 0.6797 0.4924 0.6094 0.5178 0.5382 0.5280 0.5630 0.4716 0.5173
1-shot 0.7534 0.6419 0.4828 0.6260 0.6011 0.5591 0.5801 0.5372 0.5995 0.5684
LLaMA2-chat-13B 13B 2-shot 0.7454 0.6694 0.4886 0.6345 0.5907 0.5371 0.5639 0.4633 0.5784 0.5209
4-shot 0.7471 0.7235 0.4792 0.6499 0.5750 0.5489 0.5620 0.5276 0.5555 0.5416
6-shot 0.7412 0.6911 0.5267 0.6530 0.5650 0.5527 0.5589 0.5669 0.5787 0.5728
0-Shot 0.6212 0.7128 0.6253 0.6531 0.5893 0.5669 0.5781 0.6373 0.8031 0.7202
1-Shot 0.7520 0.7172 0.6327 0.7006 0.6529 0.5968 0.6249 0.6999 0.8229 0.7614
MolecularGPT(ours) 7B 2-Shot 0.7218 0.7204 0.6338 0.6920 0.6573 0.5945 0.6259 0.7260 0.8275 0.7768
4-shot 0.7228 0.6893 0.6419 0.6847 0.6577 0.5978 0.6278 0.7168 0.8252 0.7710
6-shot 0.7181 0.6554 0.6561 0.6765 0.6629 0.5965 0.6297 0.7139 0.8289 0.7714
GCN 0.5M 0.736 0.757 0.732 0.742 0.749 0.633 0.691 0.649 0.8041 0.7266
GAT 1.0M 0.697 0.729 0.666 0.697 0.754 0.646 0.700 0.662 0.8281 0.7451
GIN 1.8M Finetuned 0.701 0.753 0.718 0.724 0.740 0.634 0.687 0.658 0.8205 0.7392
Graphormer 48M 0.7760 0.7452 0.7061 0.7424 0.7589 0.6470 0.7029 0.7015 0.8436 0.7725
Graphormer-p 48M 0.8575 0.7788 0.7480 0.7948 0.7729 0.6649 0.7189 0.7163 0.8877 0.8020
In our experimental framework, we aim to answer three primary research questions: RQ1: Can MolecularGPT effectively and robustly handle new property prediction tasks through zero- and few-shot ICL? RQ2: What is the optimal design for in-context instruction set to improve MolecularGPT’s generalization and ICL abilities during tuning? RQ3: How does the number, order, and diversity of in-context examples affect the performance of MolecularGPT?
@@ 4.1 Experimental Setup
Our baseline selection aligns with the approach used in GIMLET (Zhao et al., 2024), which can be categorized into two primary types: language models for directly inference and graph representation models totally finetuned on downstream tasks. The language models include XVPLM (Zeng et al., 2022), MoMu (Su et al., 2022), Galactica-125M (Taylor et al., 2022), Galactica-1.3B (Taylor et al., 2022), and GIMLET. And the finetuned molecular representation models comprise GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2017), GIN (Xu et al., 2019), Graphormer (Ying et al., 2021), and Graphormer-p, which is pretrained on Graphormer using datasets in GIMLET. We present the zero-shot results of these language models and the fine-tuned results of supervised models from GIMLET. Additionally, we consider general large language models: LLaMA-chat-7B and LLaMA-chat-13B as our baselines, which demonstrate ICL capabilities.
@@ 4.2 Performance Evaluation
As the results presented in Tab. 12 respectively, MolecularGPT can achieve competitive performance on classification and regression tasks under both zero-shot and few-shot settings. We answer the RQ1 with more details as follows.
The MolecularGPT demonstrates superior performance compared with other language models in zero-shot learning. In comparison to language models under zero-shot inference, our model demonstrates enhanced performance across classification and regression tasks. In terms of GIMLET, MolecularGPT surpasses it in HIV, BBBP, and CYP450 classification datasets as well as in FreeSolv and Lipo regression datasets under zero-shot condition. Compared to the LLaMA-7B and LLaMA-13B, our models exhibit a significant improvement, an average improvement of $16.6\%$ and $9.9\%$ in ROC-AUC across all classification tasks and average decrease of $5.96$ and $199.17$ in RMSE across all regression tasks correspondingly, indicating the chemical knowledge have been effectively imbued into LLaMA through our tuning.
MolecularGPT establishes a new benchmark in few-shot ICL across all tasks and outperforms the SOTA supervised models in certain conditions. When compared to the zero-shot learning, MolecularGPT demonstrates an average enhancement of $4.6\%$ in ROC-AUC across all classification tasks under one-shot condition. Compared to GIMLET, MolecularGPT exhibits an average improvement of $5.5\%$ and $5.8\%$ on classification tasks under one-shot and two-shot settings respectively. Even by one-shot ICL, MolecularGPT displays comparable performance with GCN and GIN on $3$ out of $7$ classification datasets. By two-shot ICL, MolecularGPT matches the performance of GAT on $4$ out of $7$ classification datasets. Remarkably, MolecularGPT even outperforms the highest-performing finetuned model, Graphormer-p, on BBBP dataset under two-shot condition with ROC-AUC of $0.7260$ compared to $0.7163$.
Figure 3: Standard deviation for GIMLET and MolecularGPT in response to 5 types of instructions.
Table 2: Performance on Physicalchemical regression tasks. The highlight style is the same as Tab. 1
@@
Method Type ESOL FreeSolv Lipo Avg.phy
XVPLM - - - -
MoMu 0-Shot - - - -
GIMLET 1.132 5.103 1.345 2.527
0-shot 7.227 15.912 2.329 8.489
1-shot 1.819 525.478 1.204 176.167
LLaMA2-chat-7B 2-shot 3.856 41.168 1.128 15.384
4-shot 5.940 66.593 1.112 24.548
6-shot 7.569 55.933 1.112 21.538
0-shot 281.617 321.313 2.194 201.708
1-shot 9.405 11.356 1.427 7.396
LLaMA2-chat-13B 2-shot 27.717 39.254 1.420 22.797
4-shot 643.408 9.589 1.462 218.153
6-shot 6.481 154.635 1.363 54.160
0-Shot 1.471 4.975 1.157 2.534
1-Shot 1.496 5.248 1.058 2.601
MolecularGPT(ours) 2-Shot 1.489 5.226 1.015 2.577
4-Shot 1.535 5.375 1.045 2.652
6-Shot 1.465 5.046 1.023 2.511
GCN 1.331 2.119 0.760 1.403
GAT 1.253 2.493 0.770 1.505
GIN Finetuned 1.243 2.871 0.781 1.632
Graphormer 0.901 2.210 0.740 1.284
Graphormer-p 0.804 1.850 0.675 1.110
The exceptional robustness of MolecularGPT is validated across different tasks. Given the diversity and flexibility of natural language, we aim to evaluate the robustness of MolecularGPT against various instructions. Adhering to the downstream datasets used in GIMLET, which provides five distinct types of instructions. We calculate the standard deviation of the ROC-AUC or RMSE metrics derived from these five instruction datasets. Comparative results with GIMLET is presented in Fig. 3. It is evident that our model exhibits superior robustness compared to GIMLET across most tasks. This indicts the robustness of MolecularGPT that it genuinely comprehends complex instructions and can handle a range of property prediction tasks without requiring task-specific prompt designs.

### v2

@@ Abstract
Molecular property prediction (MPP) is a fundamental and crucial task in drug discovery. However, prior methods are limited by the requirement for a large number of labeled molecules and their restricted ability to generalize for unseen and new tasks, both of which are essential for real-world applications. To address these challenges, we present MolecularGPT for few-shot MPP. From a perspective on instruction tuning, we fine-tune large language models (LLMs) based on curated molecular instructions spanning over 1000 property prediction tasks. This enables building a versatile and specialized LLM that can be adapted to novel MPP tasks without any fine-tuning through zero- and few-shot in-context learning (ICL). MolecularGPT exhibits competitive in-context reasoning capabilities across 10 downstream evaluation datasets, setting new benchmarks for few-shot molecular prediction tasks. More importantly, with just two-shot examples, MolecularGPT can outperform standard supervised graph neural network methods on 4 out of 7 datasets. It also excels state-of-the-art LLM baselines by up to 15.7% increase on classification accuracy and decrease of 17.9 on regression metrics (e.g., RMSE) under zero-shot. This study demonstrates the potential of LLMs as effective few-shot molecular property predictors. The code is available at https://github.com/NYUSHCS/MolecularGPT.
@@ 1 Introduction
  • We extensively experimented on 10 molecular property benchmarks across different scales and tasks to validate the effectiveness of MoelcularGPT. Our empirical results demonstrate that MoelcularGPT outperforms the leading LLM baselines (e.g., GIMLET and LLaMA-7b (Touvron et al., 2023)), with up to an average 15.7% improvement across all classification tasks. Additionally, with just two-shot examples, MolecularGPT surpass standard supervised GNN methods on 4 out of 7 datasets, setting new benchmarks for few-shot molecular property tasks.
@@ 2 Related work
GNNs-based MMP GNNs (Velickovic et al., 2017; Xu et al., 2019; Kipf and Welling, 2017; Stärk et al., 2022) perform MPP tasks by constructing models between molecular graphs and properties. Though these models have achieved great success (Gilmer et al., 2017; Hamilton et al., 2017; Li et al., 2022; Liu et al., 2022; Tan et al., 2023), they are implicitly trained for each task without explicit natural language instructions, limiting their ability to generalize to new tasks. While a few approaches (Guo et al., 2021; Wang et al., 2021; Schimunek et al., 2023a; Zhuang et al., 2023a) aim to extend pre-trained GNNs to new tasks, they require a large number of related tasks to pre-train a meta-learner, making them label-intensive and deviating from our generalization goals. Moreover, these supervised models rely solely on structure information, overlooking the rich textual knowledge derived from wet lab experiments.
Pretrain-finetune based molecular language models To utilize the chemical knowledge in texts, molecular language models (Liu et al., 2023a; Edwards et al., 2022; Pei et al., 2023; Zhang et al., 2023; Liu et al., 2023c) aim to integrate natural language and molecular representations for joint reasoning. These models (Su et al., 2022; Zeng et al., 2022; Liu et al., 2023b, 2024; Li et al., 2024) involve two stages: pre-training and fine-tuning. The pre-training phase primarily focuses on learning molecular representations and their associated textual descriptions through masked language modeling, contrastive learning or next token prediction. However, they still require fine-tuning on particular MPP downstream tasks, thereby limiting their generalization abilities to new tasks.
Instruction tuning based molecular language models To address this, recent efforts in molecular language modeling (Fang et al., 2023; Zhao et al., 2024) aim to explicitly align molecular graphs with their properties through instruction tuning Longpre et al. (2023). For instance, GIMLET (Zhao et al., 2024) integrates molecular graphs with instruction languages for fine-tuning LLMs. GIMLET achieves effective zero-shot ICL for new tasks but lacks few-shot ICL capability due to its generalized position embedding and decoupled attention designs. Mol-Instructions (Fang et al., 2023) is a close work to us, but it fine-tunes LLMs with only three properties tasks and neglects inter-molecular correlations, significantly limiting its zero-shot and few-shot ICL performances. In contrast, we curate a diverse instruction set covering 1000 property tasks and introduce structure-aware few-shot instructions to significantly enhance the zero-shot and few-shot reasoning capabilities of LLMs in MPP tasks. More details about these property tasks can be found in Appendix A.1
@@ 3.1 SMILES-based Molecular Instruction Tuning: A Unified Step
To improve the generalization capability of fine-tuned LLM for MPP tasks (C1), we aim to employ the well-known graph-derived linear strings (Krenn et al., 2020) of molecular graphs, SMILES (Weininger, 1988) for instruction tuning. Different from GNN encoders (Seidl et al., 2023; Zhao et al., 2024), SMILES translates molecules’ chemical structure into a string of atomic symbols and chemical bonds based on a set of rules (Qian et al., 2023). This precise translation provides a universal expression foundation for different types of molecules. Following standard instruction tuning protocol (Christofidellis et al., 2023; Zhang et al., 2023; Liu et al., 2024; Li et al., 2024), the molecular instruction set $S_{D}$ can be generated by the following prompt template $T=\{Q,I,R\}$ based on $D$, regarding as a zero-shot instruction template.
@@ 3.3 Hybrid Molecular Instruction Tuning: Better Few-Shot Learner
This observation motivates us to introduce a hybrid instruction set $S_{D}^{h}$, combining the strengths of both the zero-shot instruction template $T$ and the few-shot instruction template $T_{shot}$. Specifically, $S_{D}^{h}$ is derived from a combination of 0, 1, 2, 3, and 4-shot instruction templates. In Fig. 2, we can see that our hybrid instruction tuned models, 0&4-shot_tuning and 0-4-shot_tuning, consistently outperforms others in zero-shot and one-shot scenarios. Further details can be found in Section 4.3.
@@ 4 Experiment
Table 1: Performance over Bio-activity (Bio), Toxicity (Tox), and Pharmacokinetic (Pha) classification tasks. Highlights are the first and second best results over all 2-shot and 8-shot performances. We also report the average rank (Avg.RK) of all models under 2-shot and 8-shot separately.
Type Method Shot BACE HIV MUV Avg.Bio Tox21 ToxCast Avg.Tox BBBP CYP450 Avg.Pha Avg.RK
Few-shot Fintuned Pre-GS-Meta 2-shot 0.5213 0.5702 0.5507 0.5474 0.5976 0.5310 0.5643 0.6668 0.5105 0.5887 6.1
8-shot 0.6203 0.6128 0.5791 0.5474 0.5665 0.5208 0.5437 0.6004 0.6378 0.6191 6.4
KVPLM 2-shot 0.6380 0.6675 0.5296 0.6117 0.5336 0.5122 0.5229 0.5399 0.5960 0.5680 6.0
8-shot 0.6375 0.6293 0.4987 0.5885 0.5712 0.4892 0.5302 0.5401 0.6305 0.5853 7.0
MoMu 2-shot 0.5747 0.6435 0.4701 0.5568 0.5297 0.5164 0.5231 0.5901 0.5976 0.5939 6.7
8-shot 0.6102 0.6171 0.5804 0.6026 0.5517 0.5226 0.5372 0.5728 0.5713 0.5721 7.0
GIMLET 2-shot 0.7141 0.6794 0.6420 0.6785 0.6222 0.5903 0.6063 0.6094 0.7212 0.6653 2.4
8-shot 0.7081 0.6803 0.6487 0.6790 0.6197 0.5911 0.6054 0.6119 0.7234 0.6677 2.9
Few-shot Inference MHNfs 2-shot 0.5939 0.5835 0.5566 0.5780 0.5360 0.5370 0.5365 0.5750 0.5734 0.5742 6.0
8-shot 0.6294 0.6397 0.6764 0.6485 0.6162 0.5602 0.5892 0.5442 0.5992 0.5717 4.9
LLaMA2-7B 2-shot 0.6930 0.6587 0.5085 0.6201 0.6052 0.5010 0.5531 0.5459 0.5807 0.5633 5.9
8-shot 0.7271 0.7000 0.5161 0.6477 0.6505 0.5329 0.5917 0.5127 0.6891 0.6009 4.0
Galactica125M 2-shot 0.6298 0.6089 0.4831 0.5739 0.5902 0.5090 0.5496 0.3931 0.6961 0.5446 6.7
8-shot 0.5248 0.6474 0.5302 0.5675 0.6083 0.5035 0.5559 0.4174 0.6217 0.5196 7.0
Galactica1.3B 2-shot 0.6634 0.5295 0.5171 0.5700 0.6302 0.5566 0.5934 0.6970 0.7323 0.7147 4.0
8-shot 0.7152 0.6623 0.4977 0.6251 0.6470 0.5279 0.5875 0.7092 0.7645 0.7369 3.7
MolecularGPT 2-shot 0.7218 0.7204 0.6338 0.6920 0.6573 0.5945 0.6259 0.7260 0.8275 0.7768 1.1
8-shot 0.7331 0.6382 0.6469 0.6727 0.6565 0.5985 0.6275 0.6822 0.8228 0.7525 2.1
In our experimental framework, we aim to answer three primary research questions: RQ1: Can MolecularGPT effectively handle new property prediction tasks through few- and zero-shot ICL? RQ2: What is the optimal design for in-context instruction set to improve MolecularGPT’s generalization and ICL abilities during tuning? RQ3: How does the question prompts and retrieval in-context examples affect MolecularGPT’s performance?
@@ 4.1 Experimental Setup
We compare our MolecularGPT with 5 leading GNN-based methods: GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2017), GIN (Xu et al., 2019), Graphormer (Ying et al., 2021), and Graphormer-p (Ying et al., 2021), 5 SOTA LM-based molecular models: XVPLM (Zeng et al., 2022), MoMu (Su et al., 2022), Galactica-125M (Taylor et al., 2022), Galactica-1.3B (Taylor et al., 2022), and GIMLET (Zhao et al., 2024), and 2 strong few-shot molecular property prediction baselines: Pre-GS-Meta (Zhuang et al., 2023b) and MHNfs (Schimunek et al., 2023b), as well as the open-sourced LLM: LLaMA-chat-7B (Touvron et al., 2023)
@@ 4.2 Performance Evaluation
As the results presented in Tab. 12 respectively, MolecularGPT can achieve competitive performance on classification and regression tasks under both few-shot and zero-shot settings. We answer the RQ1 with more details as follows.
MolecularGPT establishes a new benchmark in few-shot learning. As shown in Tab. 1, our model delivers the best performance across all datasets under the few-shot setting, securing the top-1 average rank across all datasets in both the 2-shot and 8-shot scenarios, with respective ranks of 1.1 and 2.1. Most notably, unlike models in the few-shot finetuned category, such as GIMLET, which require additional tuning and parameter updates, our MolecularGPT is capable of directly inference with few-shot examples, demonstrating an average increase of 4.3% in 2-shot and 2.8% in 8-shot scenarios respectively. Comparing with strongest model in few-shot inference category, Galactica1.3B, our model shows an average increase of 7.9% in 2-shot and 3.6% in 8-shot scenarios separately, proving its strongest generalization capabilities. Additionally, MolecularGPT even surpasses supervised finetuned GNNs in 4 out of 7 classification tasks in a 2-shot setting, as shown in Tab. 1 and 2.
MolecularGPT exhibits exceptional performance in zero-shot learning. As shown in Tab. 2, during zero-shot inference, our model achieves the top-1 average rank across all datasets, outperforming the strongest baseline, GIMLET, on 5 out of 10 datasets. It demonstrates an average improvement of 6.2% on 3 classification tasks and an average decrease of 0.16 on 2 regression tasks. Comparing with our base model LLaMA, our model shows an average 15.7% improvement on classification tasks and 17.9 decrease on regression tasks. Furthermore, when compared with GNNs that have been supervised finetuned on complete training set of each dataset, our model demonstrates comparable performance on the HIV, BBBP, CYP450, and ESOL datasets, highlighting its efficacy in a zero-shot setting.
Table 2: Performance over Bio-activity, Toxicity, and Pharmacokinetic classification tasks as well as Physicalchemical regression tasks. Highlights are the first and second best results of 0-shot performances. In supervised finetuned models, we also mark the highest and lowest results. Here, "finetuned" refers to the model being trained on the complete training set.
@@
Method Type BACE HIV MUV Tox21 ToxCast BBBP CYP450 ESOL FreeSolv Lipo Avg.RK
XVPLM 0-shot Inference 0.5126 0.6120 0.6172 0.4917 0.5096 0.6020 0.5922 - - - 4.3
MoMu 0.6656 0.5026 0.6051 0.5757 0.5238 0.4981 0.5798 - - - 3.9
Galactica125M 0.4451 0.3671 0.4986 0.4964 0.5106 0.6052 0.5369 2.130 3.660 1.097 4.4
Galactica1.3B 0.5648 0.3385 0.5715 0.4946 0.5123 0.5394 0.4686 1.103 4.267 1.093 4.1
LLaMA2-7B 0.4911 0.6060 0.5554 0.5481 0.4693 0.3671 0.4198 7.227 15.912 2.329 5.6
GIMLET 0.6957 0.6624 0.6439 0.6119 0.5904 0.5939 0.7125 1.132 5.103 1.345 2.2
MolecularGPT 0.6212 0.7128 0.6253 0.5893 0.5669 0.6373 0.8031 1.471 4.975 1.157 2.1
GCN 0.736 0.757 0.732 0.749 0.633 0.649 0.8041 1.331 2.119 0.760 3.6
GAT 0.697 0.729 0.666 0.754 0.646 0.662 0.8281 1.253 2.493 0.770 4.3
GIN Finetuned 0.701 0.753 0.718 0.740 0.634 0.658 0.8205 1.243 2.871 0.781 4.0
Graphormer 0.7760 0.7452 0.7061 0.7589 0.6470 0.7015 0.8436 0.901 2.210 0.740 2.5
Graphormer-p 0.8575 0.7788 0.7480 0.7729 0.6649 0.7163 0.8877 0.804 1.850 0.675 1.0